# Co-evolution of Nuclear Star Clusters and Massive Black Holes: Extreme Mass-Ratio Inspirals

### v1

@@ Abstract
We explore extreme mass-ratio inspirals (EMRIs) in the co-evolution of massive black holes (MBHs) and nuclear star clusters (NSCs), which host diverse stellar populations across a wide range of masses. The dynamics are simulated self-consistently with GNC, which we have updated to incorporate gravitational wave orbital decay, the loss cone of a spinning MBH, and stellar evolution. Over $12$ Gyr, we investigate the evolution of the NSC with a mass-growing MBH, as well as the EMRIs of stellar black holes, neutron stars, white dwarfs, brown dwarfs (BDs), and low-mass main-sequence stars (MSs), along with tidal disruption events (TDEs) involving MSs, BDs, and post-MSs. The mass growth of the MBH contributed by TDEs is typically $\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive, Milky-Way-like, and smaller NSCs, respectively. Between $40\%$ and $70\%$ of the stellar mass is lost during stellar evolution, which dominates the mass growth of the MBH if a significant fraction of the lost mass is accreted. The evolution of EMRI rates is generally affected by the cluster’s size expansion or contraction, stellar population evolution, MBH mass growth, and the stellar initial mass function. The EMRI rates for compact objects peak at early epochs ($\lesssim 1$ Gyr) and then gradually decline over cosmic time. LISA-band ($0.1$ mHz) EMRIs involving compact objects around Milky-Way-like MBHs tend to have high eccentricities, while those around spinning MBHs preferentially occupy low-inclination (prograde) orbits. In contrast, MS- and BD-EMRIs usually have eccentricity and inclination distributions that are distinct from those of compact objects.
@@ 1 Introduction
The paper is organized as follows. In Section 2, we describe how the additional recipes are included in GNC. In Section 3, we perform simulations and investigate the evolution of the NSC, the mass growth of the MBH, and the expected properties of EMRIs in the LISA band. Discussions and conclusions are given in Section 4 and 5, respectively.
2 The method
@@
A gas reservoir formed from the mass of TDEs (Zhang & Amaro Seoane, 2025) or from mass loss during stellar evolution (see Section 2.4 and the beginning of Section 3.2) is assumed to be accreted by the MBH. The consumption rate of this gas reservoir is limited by the Eddington mass accretion rate:
@@
2.1 Including orbital decay by gravitational wave radiation
Gravitational wave emission from a stellar object around an MBH is usually considered for orbits bound to the MBH. However, since we have included the stellar potential, the orbit of a stellar object may be bound to the NSC but unbound to the MBH. If the orbit of such an object approaches the loss-cone regime, it encounters the MBH on a hyperbolic orbit. In this case, the resulting gravitational wave emission differs slightly from that of a bound orbit.
When the stellar potential is included, Keplerian orbital elements are no longer accurate, as they are defined in a two-body system. Nevertheless, since gravitational wave dissipation is most important near the pericenter, the GW emission can still be accurately estimated using the instantaneous Keplerian semi-major axis $a_{K}$ and eccentricity $e_{K}$ at the pericenter $r_{p}(E,J)$, where $r_{p}(E,J)$ is the pericenter of the orbit with energy $E$ and angular momentum $J$ (obtained by solving Equation 5 of Paper II). Since $a_{K}(1-e_{K})=r_{p}(E,J)$, we can derive the instantaneous values of $e_{K}$ and $a_{K}$:
@@ 2.1 Including orbital decay by gravitational wave radiation
For a particle bound (unbound) to the MBH, near the pericenter we instantaneously have $e_{K}<1$ and $a_{K}>0$ ($e_{K}>1$ and $a_{K}<0$).
The GW orbital decay can be added to the drift terms for orbital energy and angular momentum in Equation 3. To model EMRI events self-consistently, we calculate the GW radiation for bound or unbound orbits separately. The details are in the following two sections.
2.1.1 For orbits bound to the MBH
Let $D^{\rm GW}_{E}$ ($D^{\rm GW}_{J}$) be the drift of energy (angular momentum) due to GW radiation. For stellar orbits bound to the MBH, the orbit-averaged decay of energy and angular momentum per orbit, $\delta E^{\rm GW}$ and $\delta J^{\rm GW}$, are given by (Peters & Mathews, 1963; Peters, 1964):
$\delta E^{\rm GW}=D^{GW}_{E}P$ (5)
$=$ $-\frac{64\pi}{5}\frac{M_{\bullet}^{2}m(m+M_{\bullet})^{1/2}G^{7/2}}{r_{p}^{7/2}c^{5}}\frac{1+\frac{73}{24}e^{2}_{K}+\frac{37}{96}e^{4}_{K}}{(1+e_{K})^{7/2}},$
$\delta J^{\rm GW}=D^{GW}_{J}P$
$=$ $-\frac{64\pi}{5}\frac{M_{\bullet}^{2}mG^{3}}{r_{p}^{2}c^{5}}\frac{1+\frac{7}{8}e^{2}_{K}}{(1+e_{K})^{2}},$
2.1.2 For orbits unbound to the MBH
$\begin{gathered}\delta E^{\rm GW}=D^{GW}_{E}P=-\frac{64}{5}\frac{M_{\bullet}^{2}m(m+M_{\bullet})^{1/2}G^{7/2}}{r_{p}^{7/2}c^{5}(1+e_{K})^{7/2}}\\ \times\left[\left(\pi-\theta_{0}\right)\left(1+\frac{73}{24}e^{2}_{K}+\frac{37}{96}e^{4}_{K}\right)\right.\\ +\left.\frac{(e^{2}_{K}-1)^{1/2}}{144}\left(301+\frac{673}{2}e^{2}_{K}\right)\right],\\ \delta J^{\rm GW}=D^{GW}_{J}P=-\frac{64}{5}\frac{M_{\bullet}^{2}mG^{3}}{r_{p}^{2}c^{5}(1+e_{K})^{2}}\\ \times\left[(\pi-\theta_{0})\left(1+\frac{7}{8}e^{2}_{K}\right)+\frac{e_{K}}{8}\sin\theta_{0}(13+e^{2}_{K})\right],\end{gathered}$ (6)
respectively, where $\cos\theta_{0}=1/e_{K}$. Similarly to Section 2.1.1, here we damp the energy and angular momentum only if the particle has passed the pericenter of its orbit.
2.2 Identifying EMRIs in the simulation
@@
where $T_{\rm GW}(E,J)$ is the GW orbital decay time and $T_{\rm rlx}(E,J)$ is the two-body relaxation time for a given orbital energy $E$ and angular momentum $J$. $D_{E}^{\rm GW}$ is the energy damping rate given by either Equation 5 for bound orbits or Equation 6 for unbound orbits. $D_{JJ}^{\rm NR}$ and $D_{EE}^{\rm NR}$ are the diffusion coefficients from Equations 21 and 23 in Paper II, respectively. $C$ is a threshold parameter specifying the ratio of the two-body timescale to the GW decay timescale.
@@
Particles that fall directly into the loss cone from regimes not satisfying Equation 7 are identified as plunge events. For more details on the simulated trajectories of EMRIs and plunge events, see Section 3.1.1.
2.3 Loss cones around a spinning MBH
@@
If the MBH is spinning, the size of the ISO, $r_{\rm ISO}$, varies with the orbital inclination $i$ of the particle relative to the equatorial plane of the MBH. Following a method similar to that in Sopuerta & Yunes (2011) (see also Amaro-Seoane et al. (2013)), we derive the pericenter of the ISO, $r_{\rm ISO}$, for a particle in an inclined orbit around a spinning MBH. Generally, $r_{\rm ISO}=\mathcal{W}r_{g}$, where $\mathcal{W}$ depends on the dimensionless spin parameter $a$ and inclination $i$. For a Schwarzschild MBH ($a=0$), $\mathcal{W}=8$. If $a=1$, $\mathcal{W}$ varies from $2.12$ (prograde orbits, $i=0^{\circ}$) to $11.65$ (retrograde orbits, $i=180^{\circ}$). For more details, see Appendix A and B.
@@ 2.4 Stellar evolution
2.5 The mass-radius relation and IMF of brown dwarfs
@@
3 The dynamics of EMRIs in evolving NSCs
@@
Figure 2: Examples of evolutionary trajectories of SBHs in model M2 (see Table 1) simulated by GNC, plotted in the space of energy, angular momentum, and pericenter distance. Thin lines ending with filled red circles mark plunge events, while those ending with filled red squares mark EMRI events. Top left panel: The energy and angular momentum evolution for a maximally spinning MBH. $x=E/\sigma_{0}^{2}$ and $j=J/J_{c}(x)$ are the dimensionless energy and angular momentum (see text in Section 2 for their definitions). The loss cone size depends on the amplitude ($a$) and inclination angle (ii) of the MBH’s spin (defined in Equation A4). Top right panel: Similar to the left panel but for the evolution in the space of energy and pericenter distance. $r_{0}=3.1$ pc is a characteristic distance. Bottom panels are similar to the top panels but for the case of a Schwarzschild MBH ($a=0$). The cyan filled triangle, star, and square in each panel mark the intersection between $T_{\rm GW}=0.1T_{\rm rlx}$ and the loss cone for $a=1$, $a=0$, and a=1a=-1 (equivalent to $a=0$ with $i=180^{\circ}$), respectively.
@@
Simulations incorporating all the aforementioned recipes can provide a more comprehensive study of the evolution of NSCs, the mass growth of the MBH, and the dynamics of EMRIs. However, the coupling between these recipes substantially complicates the analysis of the simulation results. For a clear analysis of EMRI dynamics, we first adopt simplified models in Section 3.1, assuming that the NSC consists of either two or five mass components while neglecting stellar evolution. These simplified scenarios facilitate the analysis of the dynamics in later sections and comparison with many previous studies that adopted similar assumptions. We extend our analysis later in Section 3.2 to more realistic scenarios, where the NSC consists of a spectrum of mass components and includes stellar evolution.
3.1 NSCs with fixed mass components and no stellar evolution
@@
We first focus on models M2, M5, and M5_2 (Sections 3.1.13.1.2, and 3.1.3), all of which adopt a fixed MBH mass of $4\times 10^{6}M_{\odot}$, similar to the MBH in our Galactic Center (Gillessen et al., 2009). We then investigate the slightly more complicated model M5G82 (Section 3.1.4), in which the MBH mass grows from $10^{4}M_{\odot}$ by swallowing objects that fall into the loss cone (including EMRIs and plunge events) and by accreting gas released by MS-TDEs. Details of the accretion processes can be found in Section 4.2 of Paper II. For all models, the evolution and rates of TDEs are similar to those in Paper II, and thus we will not discuss them further here.
3.1.1 The formation of EMRIs
@@
3.1.2 The critical radius $a_{\rm crit}$ for different stellar objects
The intersection between $T_{\rm GW}\sim 0.1T_{\rm rlx}$ [^1] Note that here we use $C\sim 0.1$ so that the majority of EMRI events are identified within a relaxation timescale. This is different from the value $C=0.001$ described in Section 2.2, which is the critical value to identify a particle as an EMRI for an instantaneous momentum in the simulation. and the loss cone size, $j=j_{\rm lc}$, defines a critical radius $a_{\rm crit}$ (or energy $E_{\rm crit}$) below (or above) which particles are likely to become EMRIs within their relaxation time (e.g., Hopman & Alexander, 2006a). The critical radius $a_{\rm crit}$ corresponds to the radius of a circular orbit $r_{c}(E_{\rm crit})$ for the critical energy $E_{\rm crit}$. If the stellar potential is ignored, $a_{\rm crit}=GM_{\bullet}/(2E_{\rm crit})$, which reduces to the Keplerian orbital semi-major axis. The larger the value of $a_{\rm crit}$, the higher the expected rates of EMRIs.
@@ 3.1.2 The critical radius $a_{\rm crit}$ for different stellar objects
3.1.3 The evolution of the event rates
@@
These results can be understood from the discussion of $a_{\rm crit}$ in Section 3.1.2. As the $a_{\rm crit}$ of NSs, WDs, and BDs are much smaller than that of SBHs, their rates of EMRIs are orders of magnitude lower. For all compact objects, the EMRI rates are generally $\sim 3-4$ times higher around a maximally spinning MBH than around a non-spinning MBH.
@@
3.1.4 Including the MBH’s mass growth
@@
3.2 NSCs with a mass spectrum, stellar evolution, and stellar mass loss
@@
MKG91F0k $10^{9}$ $5.8\times 10^{8}$ $2.56$ $4.66$ $8.3$
Kroupa IMF
MSs ($0.01M_{\odot}-150M_{\odot}$)
0 11 $3\times 10^{7} 2.2$2.2
@@
MTG82F0.02k $1.7\times 10^{8}$ $4.1\times 10^{7}$ $1.5$ $2.7$ 2020
Top-Heavy IMF
MSs ($0.01M_{\odot}-150M_{\odot}$)
$0.02$ 11 $4.0\times 10^{6}$ $10$
@@
In this section, we extend our studies by including three additional complexities in the simulation: components with a spectrum of masses, stellar evolution processes (see details in Section 2.4), and the additional mass growth of the MBH from accreting gas released by mass loss during stellar evolution. The initial conditions of the models investigated in this section are shown in Table 2.
@@
We first provide detailed analyses of the co-evolution of the NSC and the MBH, including the evolution of stellar components (Section 3.2.1), the NSC structure (Section 3.2.2), TDEs (Section 3.2.3), and MBH mass growth (Section 3.2.4). These results are critical for understanding the formation and evolution of EMRIs, as together they establish the environmental background where EMRIs form. In Section 3.2.5, we focus specifically on models that can reproduce the Milky-Way NSC.
We then analyze the evolution of EMRI events in Section 3.2.6. The expected mass, eccentricity, and orbital inclination distributions in the LISA band will be discussed in Section 3.2.7 and 3.2.8, respectively.
3.2.1 The evolution of stellar populations

### v2

@@ Abstract
We explore extreme mass-ratio inspirals (EMRIs) in the co-evolution of massive black holes (MBHs) and nuclear star clusters (NSCs), which host diverse stellar populations across a wide range of masses. The dynamics are simulated self-consistently with GNC, which we have updated to incorporate gravitational wave orbital decay, the loss cone of a spinning MBH, and stellar evolution. Over $12$ Gyr, we investigate the evolution of the NSC with a mass-growing MBH, as well as the EMRIs of stellar black holes (SBHs), neutron stars (NSs), white dwarfs (WDs), brown dwarfs (BDs), and low-mass main-sequence stars (MSs), along with tidal disruption events (TDEs) involving MSs, BDs, and post-MSs. The mass growth of the MBH contributed by TDEs is typically $\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive, Milky-Way-like, and smaller NSCs, respectively. Between $40\%$ and $70\%$ of the stellar mass is lost during stellar evolution, which dominates the mass growth of the MBH if a significant fraction of the lost mass is accreted. The evolution of EMRI rates is generally affected by the cluster’s size expansion or contraction, stellar population evolution, MBH mass growth, and the stellar initial mass function. The EMRI rates for compact objects peak at early epochs ($\lesssim 1$ Gyr) and then gradually decline over cosmic time. LISA-band ($0.1$ mHz) EMRIs involving compact objects around Milky-Way-like MBHs tend to have high eccentricities, while those around spinning MBHs preferentially occupy low-inclination (prograde) orbits. In contrast, MS- and BD-EMRIs usually have eccentricity and inclination distributions that are distinct from those of compact objects.
@@ I Introduction
The paper is organized as follows. In Section II, we describe how the additional recipes are included in GNC. In Section III, we perform simulations and investigate the evolution of the NSC, the mass growth of the MBH, and the expected properties of EMRIs in the LISA band. Discussions and conclusions are given in Section IV and V, respectively.
II The method
@@
A gas reservoir formed from the mass of TDEs (Zhang & Amaro Seoane, 2025) or from mass loss during stellar evolution (see Section II.4 and the beginning of Section III.2) is assumed to be accreted by the MBH. The consumption rate of this gas reservoir is limited by the Eddington mass accretion rate:
@@
II.1 Including orbital decay by gravitational wave radiation
Gravitational wave emission from a stellar object around an MBH is usually considered for a Keplerian system consisting only of a particle and a point mass MBH, for both bound (Peters, 1964) and unbound (Hansen, 1972; Turner, 1977) orbits. In the case when the stellar potential is included, the Keplerian orbital energy of the particle $E_{K}=GM_{\bullet}/r-v^{2}/2=E-\phi_{\star}(r)$ is no longer conserved, as now it depends on $\phi_{\star}(r)$ at each point along the orbit. As a consequence, the Keplerian orbital elements (such as the semi-major axis $a_{K}$ and eccentricity $e_{K}$) vary along the orbit.
Since gravitational wave dissipation is most important near the pericenter, the GW emission can still be accurately estimated using the instantaneous values of $a_{K}=GM_{\bullet}/(2E_{K})$ and $e_{K}$ at the pericenter $r_{p}(E,J)$, where $r_{p}(E,J)$ is the pericenter of the orbit with energy $E$ and angular momentum $J$ (obtained by solving Equation 5 of Paper II). As a result, we have:
@@ II.1 Including orbital decay by gravitational wave radiation
$\phi_{\star}(r)$ is larger at inner regions of the cluster (see the bottom left panel of Figure 3 of Zhang & Amaro Seoane (2025)). Thus, it is possible that near the pericenter, $\phi_{\star}(r)$ is large enough that the instantaneous Keplerian orbital energy $E_{K}<0$ (thus $a_{K}<0$ and $e_{K}>1$), i.e., the orbit is instantaneously unbound to the MBH. In other words, near the pericenter the particle encounters the MBH on a hyperbolic orbit. In this case, the resulting gravitational wave emission differs slightly from that of a bound orbit.
To model EMRI events self-consistently, we calculate the GW radiation for bound or unbound orbits separately. Let $D^{\rm GW}_{E}$ and $D^{\rm GW}_{J}$ be the drifts of energy and angular momentum due to GW radiation, respectively. These GW orbital decays can be added to the drift terms for orbital energy and angular momentum in Equation 3. The details of $D^{\rm GW}_{E}$ and $D^{\rm GW}_{J}$ are shown in Appendix A.
II.2 Identifying EMRIs in the simulation
@@
where $T_{\rm GW}(E,J)$ is the GW orbital decay time and $T_{\rm rlx}(E,J)$ is the two-body relaxation time for a given orbital energy $E$ and angular momentum $J$. $D_{E}^{\rm GW}$ is the energy damping rate given by either Equation A1 for bound orbits or Equation A7 for unbound orbits. $D_{JJ}^{\rm NR}$ and $D_{EE}^{\rm NR}$ are the diffusion coefficients from Equations 21 and 23 in Paper II, respectively. $C$ is a threshold parameter specifying the ratio of the two-body timescale to the GW decay timescale.
@@
Particles that fall directly into the loss cone from regimes not satisfying Equation 5 are identified as plunge events. For more details on the simulated trajectories of EMRIs and plunge events, see Section III.1.1.
II.3 Loss cones around a spinning MBH
@@
If the MBH is spinning, the size of the ISO, $r_{\rm ISO}$, varies with the orbital inclination $i$ of the particle relative to the equatorial plane of the MBH. Following a method similar to that in Sopuerta & Yunes (2011) (see also Amaro-Seoane et al. (2013)), we derive the pericenter of the ISO, $r_{\rm ISO}$, for a particle in an inclined orbit around a spinning MBH. Generally, $r_{\rm ISO}=\mathcal{W}r_{g}$, where $\mathcal{W}$ depends on the dimensionless spin parameter $a$ and inclination $i$. For a Schwarzschild MBH ($a=0$), $\mathcal{W}=8$. If $a=1$, $\mathcal{W}$ varies from $2.12$ (prograde orbits, $i=0^{\circ}$) to $11.65$ (retrograde orbits, $i=180^{\circ}$). For more details, see Appendix B and C.
@@ II.4 Stellar evolution
II.5 The mass-radius relation and IMF of brown dwarfs
@@
III The dynamics of EMRIs in evolving NSCs
@@
Figure 2: Examples of evolutionary trajectories of SBHs in model M2 (see Table 1) simulated by GNC, plotted in the space of energy, angular momentum, and pericenter distance. Thin lines ending with filled red circles mark plunge events, while those ending with filled red squares mark EMRI events. Top left panel: The energy and angular momentum evolution for a maximally spinning MBH. $x=E/\sigma_{0}^{2}$ and $j=J/J_{c}(x)$ are the dimensionless energy and angular momentum (see text in Section II for their definitions). The loss cone size depends on the amplitude ($a$) and inclination angle (ii) of the MBH’s spin (defined in Equation B4). Top right panel: Similar to the left panel but for the evolution in the space of energy and pericenter distance. $r_{0}=3.1$ pc is a characteristic distance. Bottom panels are similar to the top panels but for the case of a Schwarzschild MBH ($a=0$). The cyan filled triangle, star, and square in each panel mark the intersection between $T_{\rm GW}=0.1T_{\rm rlx}$ and the loss cone for $a=1$, $a=0$, and a=1a=-1 (equivalent to $a=0$ with $i=180^{\circ}$), respectively.
@@
Simulations incorporating all the aforementioned recipes can provide a more comprehensive study of the evolution of NSCs, the mass growth of the MBH, and the dynamics of EMRIs. However, the coupling between these recipes substantially complicates the analysis of the simulation results. For a clear analysis of EMRI dynamics, we first adopt simplified models in Section III.1, assuming that the NSC consists of either two or five mass components while neglecting stellar evolution. These simplified scenarios facilitate the analysis of the dynamics in later sections and comparison with many previous studies that adopted similar assumptions. We extend our analysis later in Section III.2 to more realistic scenarios, where the NSC consists of a spectrum of mass components and includes stellar evolution.
III.1 NSCs with fixed mass components and no stellar evolution
@@
We first focus on models M2, M5, and M5_2 (Sections III.1.1III.1.2, and III.1.3), all of which adopt a fixed MBH mass of $4\times 10^{6}M_{\odot}$, similar to the MBH in our Galactic Center (Gillessen et al., 2009). We then investigate the slightly more complicated model M5G82 (Section III.1.4), in which the MBH mass grows from $10^{4}M_{\odot}$ by swallowing objects that fall into the loss cone (including EMRIs and plunge events) and by accreting gas released by MS-TDEs. Details of the accretion processes can be found in Section 4.2 of Paper II. For all models, the evolution and rates of TDEs are similar to those in Paper II, and thus we will not discuss them further here.
III.1.1 The formation of EMRIs
@@
III.1.2 The critical radius $a_{\rm crit}$ for different stellar objects
The intersection between $T_{\rm GW}\sim 0.1T_{\rm rlx}$ [^1] Note that here we use $C\sim 0.1$ so that the majority of EMRI events are identified within a relaxation timescale. This is different from the value $C=0.001$ described in Section II.2, which is the critical value to identify a particle as an EMRI for an instantaneous momentum in the simulation. and the loss cone size, $j=j_{\rm lc}$, defines a critical radius $a_{\rm crit}$ (or energy $E_{\rm crit}$) below (or above) which particles are likely to become EMRIs within their relaxation time (e.g., Hopman & Alexander, 2006a). The critical radius $a_{\rm crit}$ corresponds to the radius of a circular orbit $r_{c}(E_{\rm crit})$ for the critical energy $E_{\rm crit}$. If the stellar potential is ignored, $a_{\rm crit}=GM_{\bullet}/(2E_{\rm crit})$, which reduces to the Keplerian orbital semi-major axis. The larger the value of $a_{\rm crit}$, the higher the expected rates of EMRIs.
@@ III.1.2 The critical radius $a_{\rm crit}$ for different stellar objects
III.1.3 The evolution of the event rates
@@
These results can be understood from the discussion of $a_{\rm crit}$ in Section III.1.2. As the $a_{\rm crit}$ of NSs, WDs, and BDs are much smaller than that of SBHs, their rates of EMRIs are orders of magnitude lower. For all compact objects, the EMRI rates are generally $\sim 3-4$ times higher around a maximally spinning MBH than around a non-spinning MBH.
@@
III.1.4 Including the MBH’s mass growth
@@
III.2 NSCs with a mass spectrum, stellar evolution, and stellar mass loss
@@
MKG91F0k $10^{9}$ $5.8\times 10^{8}$ $2.56$ $4.66$ $8.3$ Kroupa IMF MSs ($0.01M_{\odot}-150M_{\odot}$) 0 11 $3\times 10^{7} 2.2$2.2
@@
MTG82F0.02k $1.7\times 10^{8}$ $4.1\times 10^{7}$ $1.5$ $2.7$ 2020 Top-Heavy IMF MSs ($0.01M_{\odot}-150M_{\odot}$) $0.02$ 11 $4.0\times 10^{6}$ $10$
@@
In this section, we extend our studies by including three additional complexities in the simulation: components with a spectrum of masses, stellar evolution processes (see details in Section II.4), and the additional mass growth of the MBH from accreting gas released by mass loss during stellar evolution. The initial conditions of the models investigated in this section are shown in Table 2.
@@
We first provide detailed analyses of the co-evolution of the NSC and the MBH, including the evolution of stellar components (Section III.2.1), the NSC structure (Section III.2.2), TDEs (Section III.2.3), and MBH mass growth (Section III.2.4). These results are critical for understanding the formation and evolution of EMRIs, as together they establish the environmental background where EMRIs form. In Section III.2.5, we focus specifically on models that can reproduce the Milky-Way NSC.
We then analyze the evolution of EMRI events in Section III.2.6. The expected mass, eccentricity, and orbital inclination distributions in the LISA band will be discussed in Section III.2.7 and III.2.8, respectively.
III.2.1 The evolution of stellar populations